# ST-USleepNet: A Spatial-Temporal Coupling Prominence Network for Multi-Channel Sleep Staging

### v2

@@ Abstract
Sleep staging is critical to assess sleep quality and diagnose disorders. Despite advancements in artificial intelligence enabling automated sleep staging, significant challenges remain: (1) Simultaneously extracting prominent temporal and spatial sleep features from multi-channel raw signals, including characteristic sleep waveforms and salient spatial brain networks. (2) Capturing the spatial-temporal coupling patterns essential for accurate sleep staging. To address these challenges, we propose a novel framework named ST-USleepNet, comprising a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts raw signals into a spatial-temporal graph based on signal similarity, temporal, and spatial relationships to model spatial-temporal coupling patterns. The USleepNet employs a U-shaped structure for both the temporal and spatial streams, mirroring its original use in image segmentation to isolate significant targets. Applied to raw sleep signals and graph data from the ST module, USleepNet effectively segments these inputs, simultaneously extracting prominent temporal and spatial sleep features. Testing on three datasets demonstrates that ST-USleepNet outperforms existing baselines, and model visualizations confirm its efficacy in extracting prominent sleep features and temporal-spatial coupling patterns across various sleep stages. The code is available at: https://github.com/Majy-Yuji/ST-USleepNet.git.
1Saw Swee Hock School of Public Health, National University of Singapore, Singapore
2Institute of Automation, Chinese Academy of Sciences, Beijing, China
3Institute of Data Science, National University of Singapore, Singapore
@@ 1 Introduction
Sleep disorders are associated with a wide range of health issues [1], including cardiovascular diseases [2], metabolic disorders [3], and neurodegenerative conditions [4]. Accurate diagnosis and management of these disorders heavily rely on sleep staging [5], a process that involves categorizing sleep into distinct stages based on physiological signals. This classification provides essential insights into sleep patterns and plays a critical role in identifying abnormalities.
@@
Figure 1: An illustration of brain activities during different sleep stages. (a) and (b) display characteristic sleep waveforms from different stages. (c) presents brain networks and salient subnetworks during N2 and N3. (d) illustrates different types of spatial-temporal relationships.
@@
Sleep staging is traditionally performed using polysomnography (PSG), which primarily analyzes electroencephalogram (EEG) signals, but also incorporates other physiological measurements such as electrooculogram (EOG) and electromyogram (EMG). These signals are segmented into 30-second epochs and manually classified by experts into standard stages, including wake, non-REM (N1, N2, N3) sleep stages and rapid eye movement (REM) [6]. However, this manual process is labor-intensive and challenging to scale for large datasets [7]. To improve the efficiency and accuracy of sleep staging, researchers have turned to artificial intelligence to develop automated sleep staging models [8, 9, 10]. Despite advances, there are two unresolved challenges in this field.
@@
How to simultaneously capture the characteristic sleep waveforms and salient spatial brain networks from multi-channel raw signals? Different sleep stages exhibit prominent temporal and spatial features, primarily the characteristic sleep waveforms and salient spatial brain networks as outlined by the AASM sleep standards [6]. Figures 1 (a) and (b) show characteristic sleep waveforms over time. Panel (a) features spindle waves from the N2 stage, while panel (b) displays delta waves from the N3 stage. Figure 1 (c) presents the brain networks and salient spatial subnetworks during both N2 and N3 stages, with the salient subnetworks highlighted by retaining only the green nodes. Given the crucial information contained in these prominent sleep features, many studies [11, 12, 13] have effectively extracted salient waveforms from sleep signals. However, these approaches are limited to single-channel signals, whereas PSG signals are predominantly multi-channel. The prominent sleep features in multi-channel signals are significantly more complex than those in single-channel signals. This complexity has two main sources. One arises from the characteristic sleep waveforms observed across different channels, as illustrated in Figures 1 (a) and (b), which show simultaneous waveforms in two channels. The other source comes from the formation of distinct functional connections at various sleep stages due to interactions between signals [14, 15]. These interactions result in intricate spatial patterns that further enrich the complexity of multi-channel signals. Consequently, simultaneously capturing both multi-channel characteristic sleep waveforms and salient spatial brain networks is crucial for accurate sleep staging and presents a significant challenge.
@@
How to capture the spatial-temporal coupling patterns to better understand brain activities? Signals from different brain regions exhibit complex yet crucial spatial-temporal coupling patterns as they evolve over time [16]. Figure 1 (d) illustrates the spatial-temporal relationships within the brain network over time. In the figure, each circle represents a signal channel, corresponding to an electrode placement on the brain; each arrow signifies a connection within the brain network over time. Using the green node as an example, different arrows indicate distinct types of relationships: red arrows denote temporal correlations within the same channel across different time steps, black arrows symbolize spatial dependencies at the same time step across various channels, and blue arrows represent spatial-temporal coupling patterns between different channels at different times. These spatial-temporal connections within the brain are crucial for revealing brain activities [17]. Most existing models can only capture temporal correlations [18, 19, 20, 21] or spatial dependencies [22, 23]. Therefore, capturing these spatial-temporal coupling patterns for sleep staging presents a challenge.
@@
To address the above challenges, we propose ST-USleepNet, a raw-signal-based novel spatial-temporal coupling prominence framework for sleep staging. It consists of two main components: a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts multi-channel raw signals into a spatial-temporal graph, allowing for the modeling and capturing of spatial-temporal coupling patterns. On the other hand, the USleepNet, which leverages a U-shaped structure originally developed for image segmentation [24], segments the signals and the graph constructed by the ST to obtain prominent temporal and spatial sleep features like how image segmentation isolates the main target of an image. Therefore, the USleepNet has two streams: for the temporal stream, we utilize a fully convolutional U-Net to segment characteristic sleep waveforms; for the spatial stream, we employ a graph convolutional U-Net to isolate salient spatial brain networks. The key contributions of this work include the following:
@@
  • We develop a USleepNet, comprising a temporal prominence network and a spatial prominence network to simultaneously extract multi-channel characteristic sleep waveforms and salient spatial brain networks within each sleep stage.
@@
  • We design a spatial-temporal graph construction module to model the spatial-temporal coupling patterns within a sleep segment.
@@
  • Model visualizations demonstrate that our method effectively extracts prominent sleep features and spatial-temporal coupling patterns across different sleep stages, providing a certain degree of interpretability for the method.
@@ 2 Related Work
Traditional sleep staging relied on machine learning methods like Random Forest (RF) [25] and Support Vector Machine (SVM) [26, 27], which required manually crafted features based on expert knowledge, making the process costly and inefficient. To overcome this, deep learning approaches such as convolutional neural networks (CNNs) [19, 28, 20, 21] and Recurrent Neural Networks (RNNs) [29] were introduced to automatically extract high-dimensional features. However, these methods still follow a two-stage pipeline and lack end-to-end training capabilities.
Recent efforts have focused on end-to-end architectures for processing temporal correlations in sleep staging without preliminary steps. Temporal modeling approaches include CNNs [9, 11, 30], RNNs [18], their combinations [8, 10], and Transformers [31]. For instance, U-time [9] employed a CNN-based Unet for single-channel signals, while SalientSleepNet [11] integrated a $U^{2}$-structured CNN Unet with attention to identify typical waveforms. Some studies have also focused on extracting typical waveforms to perform expert-like classification [12, 13]. Despite their advancements, these methods primarily focus on single-channel or abstract temporal features, overlooking spatial dependencies.
Spatial dependencies among brain regions are critically important [32]. GraphSleepNet [22] was the first to apply a spatial-temporal graph convolution network (ST-GCN) for sleep staging. Subsequently, researchers introduced more graph-based methods [28, 23]. Besides, Liu et al. proposed the BSTT [33] model, which uses a spatial-temporal transformer to capture spatial brain features. However, these spatial methods mainly focus on capturing global spatial features instead of the salient spatial brain networks and ignore the spatial-temporal coupling patterns.
@@ 3 Methodology
Figure 2: An illustration of ST-USleepNet. It shows the overall framework within the central gray box, while the surrounding elements detail the network layers and provide a simple example using signals from two channels.
The framework of ST-USleepNet is illustrated in Figure 2 and it consists of two primary modules. First, a spatial-temporal graph construction module is developed to capture the spatial-temporal coupling patterns, converting raw signals into a spatial-temporal graph. Next, we introduce USleepNet, which is designed to extract characteristic sleep waveforms and identify spatial salient brain networks from these signals. USleepNet is composed of two interwoven streams: the temporal stream, also called the temporal prominence network, represented by yellow blocks in Figure 2, and the spatial stream, also called the spatial prominence network, depicted by orange blocks. Each stream comprises $l$ encoder and $l$ decoder blocks, with $l=2$ exemplified in the figure. Additionally, the temporal stream integrates a segment classifier. To effectively combine the two streams, a fusion strategy is implemented.
@@ 3.1 Spatial-Temporal Graph Construction
Figure 3: An example of spatial-temporal graph construction with a 2-channel signal divided into 3 patches per channel. Each patch represents a node, and the adjacency matrix is obtained by multiplying the similarity matrix, temporal weighted mask, and positional weighted mask.
@@ 3.2.1 Temporal Prominence Network
This network consists of three logical submodules: the temporal encoder block $T_{en}$, the temporal decoder block $T_{de}$, and a segment classifier $T_{c}$. $T_{en}$ extracts multi-scale temporal features from the feature matrix $X$ of the graph $G_{s}$, outputting a compressed representation, which $T_{de}$ further processes. Last, $T_{c}$ generates a segment-wise label.
Temporal Encoder Block $T_{en}$ consists of two 1D convolution layers. The number of channels in the second convolution layer is greater than that in the first layer, since deeper networks tend to capture more complex features. Each convolution uses the same padding, followed by batch normalization. The kernel size remains consistent within the two convolution operations of each $T_{en}$, while it may vary across different $T_{en}$. After these two convolution operations, max pooling is performed to reduce the feature dimensions.
Temporal Decoder Block $T_{de}$ consists of two transposed-1D convolution layers. Each $T_{de}$ corresponds to a $T_{en}$, performing inverse operations to recover features. It unpools the compressed representation, concatenates it with the corresponding $T_{en}$ output, and applies two sequential convolutions. Kernel sizes are consistent within each $T_{de}$ but may vary across blocks, with batch normalization applied after each convolution.
Segment Classifier $T_{c}$ generates segment-wise labels. It takes embedding $H^{ch\times nC\times f_{s}t}$ as input, reshapes it to ${\hat{H}}^{ch\times C\times nf_{s}t}$, and applies a softmax function to produce point-wise labels for each time step and channel. Mean pooling combines the channels into one, and a fully connected layer outputs the segment label.
@@ 3.2.2 Spatial Prominence Network
For the spatial stream, we propose the spatial prominence network. This network follows an encoder-decoder structure and consists of two logical submodules, the spatial encoder block $S_{en}$ and the spatial decoder block $S_{de}$. $S_{en}$ takes the spatial-temporal graph $G_{s}$ as input and outputs a compressed feature subgraph, while ${S}{de}$ takes the subgraph and recovers a supergraph containing the learned spatial relationships.
Spatial Encoder Block $S_{en}$ is to extract spatial features between multiple channels and sample some important channels to form a new subgraph for high-level feature encoding. It consists of two consecutive GCN layers and a graph pooling layer.
@@
where $v$ represents the number of nodes and $h$ represents the feature dimension. Then, sort the $ic$ and select k largest values:
@@
where $idx$ is the number of selected k nodes, $\tilde{ic}^{k}$ is the weight for the top k nodes. Select the feature vectors corresponding to the top k nodes and update the adjacency matrix $\tilde{A}$ and the feature matrix $\tilde{X}$:
@@
Spatial Decoder Block $S_{de}$ is used to recover the compressed subgraph obtained by $S_{en}$ into a feature supergraph. It consists of four network layers: a convolution layer, a graph unpooling layer (gUnpool), and two graph convolution layers.
@@ 3.2.3 Spatial-Temporal Block Fusion
USleepNet comprises an alternating combination of the temporal prominence network and the spatial prominence network. It includes $l$ layers of spatial-temporal encoders (${ST}{en}$), $l$ layers of spatial-temporal decoders (${ST}{de}$), a bottom block (${ST}{b}$), and a segment classifier at the end. Specifically, each ${ST}{en}$ consists of a temporal encoder ($T_{en}$) followed by a spatial encoder ($S_{en}$), while each ${ST}{de}$ consists of a spatial decoder ($S{de}$) followed by a temporal decoder ($T_{de}$). The bottom block (${ST}{b}$) is similar to ${ST}{en}$ but does not include any downsampling. In the encoder phase, it is essential to process with $T_{en}$ first, as PSG data typically has high dimensions and contains significant noise. Convolution and pooling operations not only reduce dimensions but also remove noise, thereby extracting more representative features. These lower-dimensional features significantly reduce the computational complexity of subsequent graph processing.

### v3

@@ Abstract
Sleep staging is critical to assess sleep quality and diagnose disorders. Despite advancements in artificial intelligence enabling automated sleep staging, significant challenges remain: (1) Simultaneously extracting prominent temporal and spatial sleep features from multi-channel raw signals, including characteristic sleep waveforms and salient spatial brain networks. (2) Capturing the spatial-temporal coupling patterns essential for accurate sleep staging. To address these challenges, we propose a novel framework named ST-USleepNet, comprising a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). The ST module converts raw signals into a spatial-temporal graph based on signal similarity, temporal, and spatial relationships to model spatial-temporal coupling patterns. The USleepNet employs a U-shaped structure for both the temporal and spatial streams, mirroring its original use in image segmentation to isolate significant targets. Applied to raw sleep signals and graph data from the ST module, USleepNet effectively segments these inputs, simultaneously extracting prominent temporal and spatial sleep features. Testing on three datasets demonstrates that ST-USleepNet outperforms existing baselines, and model visualizations confirm its efficacy in extracting prominent sleep features and temporal-spatial coupling patterns across various sleep stages. The code is available at https://github.com/Majy-Yuji/ST-USleepNet.
@@ 1 Introduction
Sleep disorders are associated with a wide range of health issues Mahowald and Schenck (2005), including cardiovascular diseases Redline et al. (2023), metabolic disorders Gileles-Hillel et al. (2016), and neurodegenerative conditions Abbott and Videnovic (2016). Accurate diagnosis and management of these disorders heavily rely on sleep staging Malhotra and White (2002), a process that involves categorizing sleep into distinct stages based on physiological signals. This classification provides essential insights into sleep patterns and plays a critical role in identifying abnormalities.
@@
Figure 1: Illustration of brain activities across sleep stages: (a–b) display characteristic sleep waveforms from different stages. (c) presents brain networks and salient subnetworks during N2 and N3. (d) illustrates different types of spatial-temporal relationships.
@@
Sleep staging is traditionally performed using polysomnography (PSG), which primarily analyzes electroencephalogram (EEG) signals, but also incorporates other physiological measurements such as electrooculogram (EOG) and electromyogram (EMG). These signals are segmented into 30-second epochs and manually classified by experts into standard stages, including wake, non-REM (N1, N2, N3) sleep stages and rapid eye movement (REM) Berry et al. (2012). However, this manual process is labor-intensive and challenging to scale for large datasets Danker-Hopfe et al. (2009). To improve the efficiency and accuracy of sleep staging, researchers have turned to artificial intelligence to develop automated sleep staging models Supratak et al. (2017); Perslev et al. (2019); Supratak and Guo (2020). Despite advances, there are two unresolved challenges in this field.
@@
How to simultaneously capture the characteristic sleep waveforms and salient spatial brain networks from multi-channel raw signals? Different sleep stages exhibit prominent temporal and spatial features, primarily the characteristic sleep waveforms and salient spatial brain networks as outlined by the AASM sleep standards Berry et al. (2012). Figures 1 (a) and (b) show characteristic sleep waveforms over time. Panel (a) features spindle waves from the N2 stage, while panel (b) displays delta waves from the N3 stage. Figure 1 (c) presents the brain networks and salient spatial subnetworks during both N2 and N3 stages, with the salient subnetworks highlighted by retaining only the green nodes. Given the crucial information contained in these prominent sleep features, many studies Jia et al. (2021b); Pei et al. (2024b); Wang et al. (2024a) have effectively extracted salient waveforms from sleep signals. However, these approaches are limited to single-channel signals, whereas PSG signals are predominantly multi-channel. The prominent sleep features in multi-channel signals are significantly more complex than those in single-channel signals. This complexity has two main sources. One arises from the characteristic sleep waveforms observed across different channels, as illustrated in Figures 1 (a) and (b), which show simultaneous waveforms in two channels. The other source comes from the formation of distinct functional connections at various sleep stages due to interactions between signals Stevner et al. (2019); Parente and Colosimo (2020). These interactions result in intricate spatial patterns that further enrich the complexity of multi-channel signals. Consequently, simultaneously capturing both multi-channel characteristic sleep waveforms and salient spatial brain networks is crucial for accurate sleep staging and presents a significant challenge.
@@
How to capture the spatial-temporal coupling patterns to better understand brain activities? Signals from different brain regions exhibit complex yet crucial spatial-temporal coupling patterns as they evolve over time Hövel et al. (2020). Figure 1 (d) illustrates the spatial-temporal relationships within the brain network over time. In the figure, each circle represents a signal channel, corresponding to an electrode placement on the brain; each arrow signifies a connection within the brain network over time. Using the green node as an example, different arrows indicate distinct types of relationships: red arrows denote temporal correlations within the same channel across different time steps, black arrows symbolize spatial dependencies at the same time step across various channels, and blue arrows represent spatial-temporal coupling patterns between different channels at different times. These spatial-temporal connections within the brain are crucial for revealing brain activities Pang et al. (2023). Most existing models can only capture temporal correlations Phan et al. (2018); Eldele et al. (2021); Ji et al. (2023, 2024) or spatial dependencies Jia et al. (2020); Ji et al. (2022). Therefore, capturing these spatial-temporal coupling patterns for sleep staging presents a challenge.
@@
To address the above challenges, we propose ST-USleepNet, a raw-signal-based novel spatial-temporal coupling prominence framework for sleep staging. It consists of two main components: a spatial-temporal graph construction module (ST) and a U-shaped sleep network (USleepNet). Inspired by studies from other domains Song et al. (2020); Gong et al. (2023), the ST module converts multi-channel raw signals into a spatial-temporal graph, allowing for the modeling and capturing of spatial-temporal coupling patterns. On the other hand, the USleepNet, which leverages a U-shaped structure originally developed for image segmentation Ronneberger et al. (2015), segments the signals and the graph constructed by the ST to obtain prominent temporal and spatial sleep features like how image segmentation isolates the main target of an image. Therefore, the USleepNet has two streams: for the temporal stream, we utilize a fully convolutional U-Net to segment characteristic sleep waveforms; for the spatial stream, we employ a graph convolutional U-Net to isolate salient spatial brain networks. The key contributions of this work include the following:
@@
  • We develop the USleepNet module, comprising a temporal prominence network and a spatial prominence network to simultaneously extract multi-channel characteristic sleep waveforms and salient spatial brain networks.
@@
  • We design a spatial-temporal graph construction module to model the spatial-temporal coupling patterns.
@@
  • Visualization Studies demonstrate that our method effectively extracts prominent sleep features and spatial-temporal coupling patterns across different sleep stages, providing a certain degree of interpretability.
@@ 2 Related Work
Figure 2: ST-USleepNet overview: central box shows the framework; surrounding parts illustrate layers and a two-channel example.
Traditional sleep staging relied on machine learning methods like Random Forest (RF) Memar and Faradji (2018) and Support Vector Machine (SVM) Lv et al. (2016); Alickovic and Subasi (2018), which required manually crafted features based on expert knowledge, making the process costly and inefficient. To overcome this, deep learning approaches such as convolutional neural networks (CNNs) Eldele et al. (2021); Jia et al. (2021a); Ji et al. (2023, 2024) and Recurrent Neural Networks (RNNs) Pei et al. (2024a) were introduced to automatically extract high-dimensional features. However, these methods still follow a two-stage pipeline and lack end-to-end training capabilities.
Recent efforts have focused on end-to-end architectures for processing temporal correlations in sleep staging without preliminary steps. Temporal modeling approaches include CNNs Perslev et al. (2019); Jia et al. (2021b); Perslev et al. (2021), RNNs Phan et al. (2018), their combinations Supratak et al. (2017); Supratak and Guo (2020), and Transformers Phan et al. (2022). For instance, U-time Perslev et al. (2019) employed a CNN-based Unet for single-channel signals, while SalientSleepNet Jia et al. (2021b) integrated a $U^{2}$-structured CNN Unet with attention to identify typical waveforms. Some studies have also focused on extracting typical waveforms to perform expert-like classification Pei et al. (2024b); Wang et al. (2024a). Despite their advancements, these methods primarily focus on single-channel or abstract temporal features, overlooking spatial dependencies.
Spatial dependencies among brain regions are critically important Sakkalis (2011). GraphSleepNet Jia et al. (2020) was the first to apply a spatial-temporal graph convolution network (ST-GCN) for sleep staging. Subsequently, researchers introduced more graph-based methods Jia et al. (2021a); Ji et al. (2022). Besides, Liu et al. proposed the BSTT Liu and Jia (2023) model, which uses a spatial-temporal transformer to capture spatial brain features. However, these spatial methods mainly focus on capturing global spatial features instead of the salient spatial brain networks and ignore the spatial-temporal coupling patterns.
@@ 3 Methodology
The framework of ST-USleepNet is illustrated in Figure 2, consisting of two modules. First, a spatial-temporal graph construction module is developed to capture the spatial-temporal coupling patterns, converting raw signals into a spatial-temporal graph. Next, we introduce USleepNet to extract characteristic sleep waveforms and identify spatial salient brain networks from these signals. USleepNet is composed of two interwoven streams: the temporal stream, also called the temporal prominence network, represented by yellow blocks in Figure 2, and the spatial stream, also called the spatial prominence network, depicted by orange blocks. Each stream comprises $l$ encoder and $l$ decoder blocks, with $l=2$ exemplified in the figure. Additionally, the temporal stream integrates a segment classifier. To effectively combine the two streams, a fusion strategy is implemented.
@@ 3.1 Spatial-Temporal Graph Construction
Figure 3: Spatial-temporal graph from a 2-channel signal (3 patches/channel); nodes are patches, and adjacency is computed via similarity, temporal, and positional masks.
@@ 3.2.1 Temporal Prominence Network
This network consists of three logical submodules: the temporal encoder block $T_{en}$, the temporal decoder block $T_{de}$, and a segment classifier $T_{c}$. $T_{en}$ extracts multi-scale temporal features from the feature matrix $X$ of the graph $G_{s}$, outputting a compressed representation, which $T_{de}$ further processes. Lastly, $T_{c}$ generates a segment-wise label.
Temporal Encoder Block
$T_{en}$ consists of two 1D convolution layers. The number of channels in the second convolution layer is greater than that in the first layer, since deeper networks tend to capture more complex features. Each convolution uses the same padding, followed by batch normalization. The kernel size remains consistent within the two convolution operations of each $T_{en}$, while it may vary across different $T_{en}$. After these two convolution operations, max pooling is performed to reduce the feature dimensions.
Temporal Decoder Block
$T_{de}$ consists of two transposed-1D convolution layers. Each $T_{de}$ corresponds to a $T_{en}$, performing inverse operations to recover features. It unpools the representation, concatenates it with the corresponding $T_{en}$ output, and applies two sequential convolutions. Kernel sizes are consistent within each $T_{de}$ but may vary across blocks, with batch normalization applied after each convolution.
Segment Classifier
$T_{c}$ generates segment-wise labels. It takes embedding $H^{ch\times nC\times f_{s}t}$ as input, reshapes it to ${\hat{H}}^{ch\times C\times nf_{s}t}$, and applies a softmax function to produce point-wise labels for each time step and channel. Mean pooling combines the channels into one, and a fully connected layer outputs the segment label.
@@ 3.2.2 Spatial Prominence Network
For the spatial stream, inspired by Graph U-Net Gao and Ji (2019), we propose the spatial prominence network. This network follows an encoder-decoder structure and consists of two logical submodules, the spatial encoder block $S_{en}$ and the spatial decoder block $S_{de}$. $S_{en}$ takes the spatial-temporal graph $G_{s}$ as input and outputs a compressed feature subgraph, while ${S}{de}$ takes the subgraph and recovers a supergraph containing the learned spatial relationships.
Spatial Encoder Block
$S_{en}$ is to extract spatial features between multiple channels and sample important channels to form a new subgraph for high-level feature encoding. It consists of two GCN layers and a graph pooling layer.
@@
where $v$ represents the number of nodes, $h$ represents the feature dimension. Then, sort the $ic$ and select k largest values:
@@
where $idx$ is the number of selected top-$k$ nodes, and $\tilde{ic}^{k}$ is their weights. Feature vectors of the top-$k$ nodes are used to update the adjacency matrix $\tilde{A}$ and feature matrix $\tilde{X}$:
@@
Spatial Decoder Block
$S_{de}$ is used to recover the compressed subgraph obtained by $S_{en}$ into a feature supergraph. It consists of four network layers: a convolution layer, a graph unpooling layer (gUnpool), and two graph convolution layers.
@@ 3.2.3 Spatial-Temporal Block Fusion
USleepNet comprises an alternating combination of the temporal prominence network and the spatial prominence network. It includes $l$ layers of spatial-temporal encoders (${ST}{en}$), $l$ layers of spatial-temporal decoders (${ST}{de}$), a bottom block (${ST}{b}$), and a segment classifier at the end. Specifically, each ${ST}{en}$ consists of a temporal encoder ($T_{en}$) followed by a spatial encoder ($S_{en}$), while each ${ST}{de}$ consists of a spatial decoder ($S{de}$) followed by a temporal decoder ($T_{de}$). The bottom block (${ST}{b}$) is similar to ${ST}{en}$ but does not include any downsampling. In the encoder phase, it is essential to process with $T_{en}$ first, as PSG data typically has high resolutions and contains noise. Convolution and pooling operations not only reduce dimensions but also remove noise, thereby extracting more representative features. These lower-dimensional features significantly reduce the computational complexity of subsequent graph processing.